OpenStudy (anonymous):

Solve for x. 5x - 1 = 26 a. x = 27/5 b. x = 5 c. x = -5

OpenStudy (kirstin012596):

5x-1=26 +1 +1 5x=27 (divide by 5 on both sides to get x alone) 27/5=5.4 x=5.4

OpenStudy (anonymous):

5.4